Opened 5 years ago

Closed 3 years ago

Last modified 22 months ago

#560 closed flight processing (fixed)

MA14/10, flight day 316/2014, Berkelah Forest Reserve

Reported by: dap Owned by:
Priority: whenever Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~arsf/arsf_data/2014/flight_data/malaysia/MA14_10-2014_316_Berkelah

Data arrived from ARSF via network transfer on 05/01/2015.

Scientific objective: To increase knowledge of disturbance and regeneration in tropical forests to support sustainable forest management.

Priority: Unknown

PI: Alan Blackburn

PI is happy to receive data by network transfer, USB hard disk or DVD.

Sensors:

  • Fenix (requested)
  • LiDAR (requested)
  • FW LiDAR (requested)

RCD images were requested but could not be collected due to an instrument failure.

Change History (27)

comment:1 Changed 5 years ago by dac

Received email from Nurul Ain (gmail - nurulainzaki) a PhD student co-supervised by Alan Blackburn 13/03/2015 asking to be notified on processing status. Requested data be made available via FTP.

Version 0, edited 5 years ago by dac (next)

comment:2 Changed 4 years ago by dac

  • Description modified (diff)

comment:3 Changed 4 years ago by stgo

Navigation Processing
Have had to convert the basestation using teqc
start time: 2014 Nov 11 22:46:56.780
end time: 2014 Nov 12 02:36:00.020

Basestation position

lat 3 07 31.68823
long 101 33 43.39241
height 18.976

comment:4 Changed 4 years ago by stgo

Lidar

Starting processing

comment:5 Changed 4 years ago by asm

Hyperspectral
Starting processing

comment:6 Changed 4 years ago by asm

Found SCTs values:

Flightline FENIX
1 0.90
2 0.97
3 0.97
4 0.90
5 -0.03
6 0.90
7 0.94
8 0.94
9 0.90
10 0.90
11 0.87
12 0.89
13 -0.05
14 1.00
15 0.82
16 1.00
17 0.72
18 0.98
19 0.95
20 1.05
21 0.92
Last edited 4 years ago by asm (previous) (diff)

comment:7 Changed 4 years ago by asm

Hyperspectral Proccessing
Ready for delivery check.

comment:8 Changed 4 years ago by stgo

Lidar Processing

Full waveform processing complete, ready for delivery check.

Pitch and roll values are noted below:

Pitch-0.00265
Roll-0.00635

comment:9 Changed 4 years ago by gej

Lidar Deliver Check
All looks good, although mosaic screenshot is difficult to see in readme due to flight line 30 being so far away(this is mentioned in read me so should be fine)

NOTE: First delivery check so should be checked by someone else

comment:10 Changed 4 years ago by dac

LiDAR DC

Starting second check.

comment:11 Changed 4 years ago by dac

LiDAR DC

Looks OK.

  • las1.0 folder should be las1.2 - Have renamed (we really should update the make delivery script)
  • The intensity image looks silly, suggest splitting into two figures in the readme. Would also need to do with DEM so they match. Suggest using linked displays in ENVI / TuiView and taking screengrabs from these.

comment:12 Changed 4 years ago by stgo

Lidar Processing

Have moved the new readme into the delivery folder and also copied the new screenshots focusing on the data areas in. Ready for re DC.

Last edited 4 years ago by stgo (previous) (diff)

comment:13 Changed 4 years ago by gej

Lidar DC

New Dems and screenshots look fine aswell as the ReadMe.

Second check?

comment:14 Changed 4 years ago by dac

LiDAR DC

Looks fine, ready to delivery.

comment:15 Changed 4 years ago by gej

Lidar delivered to PI via FTP

comment:16 Changed 4 years ago by dac

LiDAR Delivery

Emailed Nurul Ain informing them project was available via FTP and to get login details from Alan Blackburn. Didn't send FTP details directly as wasn't listed on application and haven't been informed by PI we should make data available.

Last edited 4 years ago by dac (previous) (diff)

comment:17 Changed 4 years ago by dac

LiDAR

The script to produce KMLs was giving an error about points being outside the DEM. Rechecked and the DEM is smaller than the hyperspectral coverage (by ~ 0.8 km), this is due to the two lines being so far apart and having to use the lidar extent to patch to the ASTER with rather than using the bounding box of the navigation data.

Have reprocessed using a buffer of 4 km. Emailed PI (Alan Blackburn)to inform him of the update and have placed files on the FTP server.

comment:18 Changed 4 years ago by stgo

Fenix reprocessing

Ready for delivery check.

This delivery needs to be thoroughly checked for:

  • correct lines included
  • correct linenames from logsheet
  • data quality comments
  • spectra shifts

comment:19 Changed 4 years ago by gej

Fenix Delivery Check
Have moved logsheet into delivery.

Screenshots need renaming according to proj_tidy.

Data quality report needs updating to 2015 version

Have checked spectra and looks okay to me.

Last edited 4 years ago by gej (previous) (diff)

comment:20 Changed 4 years ago by stgo

Fenix DC

renamed screenshots.

Have updated data quality report and edited projxml to reflect new year.

comment:21 Changed 4 years ago by gej

Fenix DC

Screenshots now correct. data quality report and prjxml correct.

Spectra needs checking by either lah or dac

comment:22 Changed 4 years ago by dac

Fenix Spectra Check

Spectra look too low compared to Py6S. Also checked against 2014/169 and much lower. Need to investigate further before delivering.

comment:23 Changed 4 years ago by stgo

Fenix delivery

Updated the readme file with data quality comment on low spectra

comment:24 Changed 4 years ago by dac

Fenix Delivery Check

Have checked spectra again. Apart from magnitude being lower than expected are fine, shape as expected, no discontinuity between VNIR and SWIR. Readme is fine. Add note to delivery email, can investigate further with PI if needed.

Ready to deliver.

comment:25 Changed 4 years ago by gej

Fenix Delivered to Dr. Alan Blackburn via ftp on arsf16 on December 9th 2015

comment:26 Changed 3 years ago by lah

Archiving
Checked and uploaded to NEODC 20/06/16.

Last edited 3 years ago by lah (previous) (diff)

comment:27 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Raw and processed data archived at NEODC. Closing ticket,

Note: See TracTickets for help on using tickets.